The attractive force between two equal but opposite charges is F newtons when the charges are 6 cm apart. What should be their distance apart to increase the attractive force to 6F newtons?

1
Expert's answer
2022-03-21T12:47:45-0400

We know that

"F=\frac{kq_1q_2}{r^2}"

Force inverse square proportional to distance

"\frac{F_1}{F_2}=\frac{r_2^2}{r_1^2}"

"\frac{F}{6F}=\frac{r_2^2}{0.06^2}"

"\frac{36\times10^{-4}}{6}=r_2^2"

"r_2=\sqrt{6\times10^{-4}}=0.0244m\\r_2=2.44cm"